Individual retirement accounts (IRA) allow some taxpayers to set aside kshs.160000 a year for retirement, and the interest earned on these accounts is exempt from taxation. If an individual starts selling aside money in an IRA early in her working life, its value at retirement can be substantially higher than the amount actually put in. Wanjiku sets aside kshs.160, 000 a year and starting when she is 25 years for an expected retirement age of 65, and expects to make 8% p.a return on investments. Calculate the total expected value when she retires. What if the income is taxed at 40% per year
1
Expert's answer
2019-04-03T09:03:55-0400

Using annuity formula:

"FV = \frac{160,000*((1 + 0.08)^{40} - 1)}{0.08} = 41,449,043."

If 40% tax is deducted, then:

"FV = \frac{160,000*((1 + 0.08*(1 - 0.4))^{40} - 1)}{(0.08*0.6)} = 18,410,196.6."